OpenStudy (anonymous):

HELP?!?! Suppose you divide a polynomial by a binomial. How do you know if the binomial is a factor of the polynomial? Create a sample problem that has a binomial which IS a factor of the polynomial being divided, and another problem that has a binomial which is NOT a factor of the polynomial being divided.

OpenStudy (anonymous):

Do it backward. Start with a binomial; let's pick x + 1. Now multiply that by another binomial; let's pick x - 2. Multiply those together using FOIL: (x + 1)(x - 2) = x^2 - 2x + 1x - 2 = x^2 -x - 2 Now we have a polynomial, and we already know that a binomial (x +1) is a factor of it. So our sample problem could be (x^2 - x - 2) / (x + 1) Done!

OpenStudy (anonymous):

thank you ace(:

OpenStudy (anonymous):

no problem :D